two years ago, sue was half as old as pat. eight years from now, sue will be two-thirds as old as pat. how old is each of them now?
Answer by stanbon(75887)

Two years ago, sue was half as old as pat. eight years from now, sue will be two-thirds as old as pat. how old is each of them now?
------------
Equations:
(s-2) = (1/2)(p-2)
(s+8) = (2/3)(p+8)
------
Rearrange:
2s-4 = p-2
3s+24= 2p+16
-----
Rearrange:
2s-p = 2
3s-2p = -8
----
Multiply thru 1st eq. by 2 to get:
4s-2p = 4
3s-2p = -8
---
Subtract 2nd from 1st to get:
s = 12 (Sue's age now)
---
Since 4s-2p = 4, substitute for "s" and solve for "p":
4*12-2p = 4
48-2p = 4
2p = 44
p = 22 (Pat's age now)
==========
Cheers,
Stan H.
=========